# spacing=nonfrench

### v1

@@ 5.1. LibLPM
LibLPM (Rasiuckevicius, 2022) is an open-source Longest Prefix Match (LPM) library written in C with built-in support for IPv4 and IPv6 addresses. LPM is an important operation that is used in packet routing for finding the best and most accurate matches in routing tables for incoming packets.
To integrate LibLPM with Iridescent, we create two different specializations - (i) LibLPM-FP, and (ii) LibLPM-NI.
LibLPM-FP specializes the lookup function to add a fast-path specialization point. The target is the lookup function where the input-output pairs of the lookup function executions are captured.
@@
LibLPM-NI creates a code-generation specialization point in the LPM library. The code-generation specialization point generates a new version of the lookup function which generates a nested-if-else tree of checks consisting of prefix match checks for the incoming address for each lpm entry. The nested-if-else tree starts at the least specific matching rule (based on prefix length) and progressively checks for the most specific matching rule until it can’t find one anymore. By performing this code generation specialization, the lookup function can embed the prefix rules directly into the codebase allowing for more optimized checks.
@@
We also create an additional specialization, LibLPM-NI-FP, which combines the previous two specializations by adding a fast-path specialization point in the generated nested-if specialized lookup function.
5.2. TAS
@@
TAS (Kaufmann et al., 2019), TCP acceleration as a service, is a lightweight software TCP network fast-path that is optimized for common-case client-server RPCs. TAS executes common-case TCP operations in an isolated fast path that uses DPDK (DPDK Project, 2022), while handling corner cases in a slow path.
@@
To integrate TAS with Iridescent, we convert the BATCH_SIZE as a runtime constant specialization point. In TAS, the BATCH_ SIZE variable is used in three different scenarios: (i) to determine the number of packets to be read from the NIC, (ii) to determine the number of packets read from the application queues, and (iii) to determine the number of packets read from the queue manager. For more fine-grained control, we convert each of these usage instances of the BATCH_SIZE as three separate runtime constant specialization points - (i) rx_batch, (ii) queues_batch, and (iii) qman_batch.
5.3. FastClick
@@
FastClick (Barbette et al., 2015) is an extended version of the Click Modular Router (Kohler et al., 2000) with improved Netmap support and DPDK support for running the modular router in userspace. A FastClick (or Click) router is assembled from individual packet processing modules called elements. Each element implements simple router functions such as packet classification, scheduling, routing, and interacting with network devices. Users can configure the router with different element pipelines to achieve different router behaviors.
@@
To integrate FastClick with Iridescent, we modify the LinearIPLookup element of the FastClick to create a Iridescent-enabled LinearIPLookup element. The original element uses a linear search algorithm to find the best matching route for an incoming packet based on the routine table. We modify the packet-processing function of the LinearIPLookup element to add a fast-path specialization point. The target of this specialization point is the internal lookup function of the LinearIPLookup element.
5.4. Network Functions
@@
Network Functions is a suite of the network functions that include DPDK and ebpf implementations of common network functions such as a NAT, Router, policer, among others. These functions have been developed and used for evaluation by various research projects such as Pix (Iyer et al., 2022b) and Vigor (Zaostrovnykh et al., 2019). We extract the implementations from the Pix artifact (Iyer et al., 2022a).
@@
6. Evaluation
In this section, we evaluate how well Iridescent improves performance of systems by enabling online specializations. We showcase how we can use online specialization with Iridescent for three use-cases: (i) enabling compile-time optimizations at runtime; (ii) enabling incremental specializations at runtime; (iii) design exploration at runtime. We answer the following questions:
  • Do Iridescent-enabled compile-time optimizations at runtime improve performance? (§ 6.2)
  • Can Iridescent find an optimal design through exploration at runtime under dynamic conditions? (§ 6.3)
  • What is the cost of using Iridescent? (§ 6.4)
6.1. Experimental Setup
Testbed. For our experiments, unless otherwise stated, we configure two identical machines as client and server. They are directly connected with a pair of 100 Gbps Mellanox ConnectX-5 Ethernet adapters. Both machines have two Intel Xeon Gold 6152 processors at 2.1 GHz, each with 22 cores for a total of 44 cores and 187 GB of RAM per machine. We run Linux kernel 5.15 with Debian 11.
6.2. Compiler Optimizations at Runtime
@@
Table 3. Impact of turning $s$ as compile-time constant at runtime for N=64 for different hardwares (Table 1)
Figure 3. Average number of cycles required to perform one lookup for different number of entries in the lookup table for two different workloads.
Compile-time optimizations often produce more performant code. These compile time optimizations include constant propagation, loop unrolling, dead-code elimination, and use of vectorization instructions. With Iridescent, we can enable compile time optimizations at runtime.
Iridescent enables Cascading Compiler Optimizations. First, we show the potential benefits of enabling compile-time optimizations at runtime for the MMulBlockBench microbenchmark shown in Table 1. Specifically, for the workload $N=64$, we compare the performance of keeping the optimal block size,$B$, as a runtime parameter to that of converting it into a constant at runtime using Iridescent. For both configurations, we execute the function for a fixed number of times and measure the amount of cycles spent for executing 1 execution of the function. Table 3 shows the benefit for converting $B$ into a compile-time constant for different hardwares. For each hardware, we get at least 50% reduction in consumed cycles, and greater than 240% reduction in consumed cycles for four of the five operating conditions. This improvement is a direct impact of Iridescent’s specialization allowing different compiler optimizations to cascade and combine together to produce a more efficient version of the code. In this specific case, the specialization of $B$ allows the JIT compiler to first easily unroll the loops in the matrix multiplication function and then replace multiple instructions with more optimized vector instructions.
Figure 4. Change in overall throughput of the FastPath router with Iridescent-specialized LinearIPLookup element for different sizes of the overall throughput.
Figure 5. Change in overall throughput of the FastPath router with Iridescent-specialized LinearIPLookup element for different fast-path hit rates.
@@
Iridescent enables custom specializations. Next, we show the potential benefits of enabling custom compile-time optimizations for LibLPM. For this experiment, we set up eleven different configurations resulting from a combination of five different LPM table sizes and two different workloads - Workload A and Workload B. In Workload A, the incoming IP address is an IP address that matches a very specific prefix entry in the routing table. In Workload B, the incoming IP address is an IP address that only matches the LPM prefix entry of prefix length 0. Thus, Workload A and Workload B cover the best and worst cases respectively for the LPM lookup function. We execute these workloads on seven different specializations - (i) Baseline: no Iridescent-enabled specialization, (ii) Instrumented: Iridescent-enabled specialization which captures the most popular inputs (with a 10% sampling rate); information collected by this configuration is used by Iridescent for the fast path specializations, (iii) LibLPM-FP-A: fast-path specialization of size 1 specialized for Workload A, (iv) LibLPM-FP-B: fast-path specialization of size 1 specialized for Workload B, (v) LibLPM-NI: the nested-if code generation specialization, (vi) LibLPM-NI-FP-A: Combination of LibLPM-NI and LibLPM-FP-A, and (vii) LibLPM-NI-FP-B: Combination of LibLPM-NI and LibLPM-FP-B. Figure 10 shows the cycles required for completing one execution on average of the lpm lookup function for all the different combinations of configurations and specializations. For Workload A, all three of LibLPM-FP-A, LibLPM-NI, LibLPM-NI-FP-A specializations achieve upto $6\times$ reduction in cycles per execution. However, unlike the LibLPM-FP and LibLPM-NI-FP-A specialization, this LibLPM-NI specialization does not require an additional instrumentation phase to sample executions to find the optimal candidates for the fast path. LibLPM-NI improves performance across the board for all workloads by upto $6\times$. Despite this performance boost, LibLPM-NI is not the best specialization for Workload B. For Workload B, the best optimizations are the *-FP-B specializations as they get upto a $30\times$ performance boost.
6.3. Runtime Design Exploration
Figure 6. Automatic Exploration and Specialization of different configurations for TAS
@@ 6.3. Runtime Design Exploration
Iridescent enables runtime design exploration. We show runtime design exploration for TAS (Kaufmann et al., 2019). We setup our experiment with a typical server-client setting, with the server providing an echo service. The client is multi-threaded and executes an open-loop workload of 64 byte packets. Both the server and client are enabled with TAS. In our experiment, we modify the server-side TAS and measure the throughput of the server side TAS as millions of packets processed per second. With Iridescent we explore different values of the rx_batch in the TAS source code on the server side. Figure 6 shows how Iridescent can automatically select the best-configuration on the server side by automatically exploring the different values of the batch size for different locations in the source code.
Figure 7. Automatic Exploration and Specialization of MMulBlockBench for two different workloads
@@
Iridescent can automatically adapt to changing workloads. To show that Iridescent can automatically adapt to workloads, we first use the MMulBlockBench microbenchmark. We execute a Iridescent-enabled version of the function and a non-Iridescent enabled baseline version. In the non-baseline version, the block size $s$ is a runtime parameter. In the Iridescent-enabled version, Iridescent explores different values of $s$ by specializing the value in the code and converting $s$ into a compile-time constant. We execute each version with two different workloads in succession, each lasting one minute. We measure the overall throughput (executions/s) for both configurations. Figure 7 shows how Iridescent can easily explore the different block sizes and find a performant configuration as compared to the baseline. Moreover, Iridescent can automatically detect the workload change based on the drop in throughput and restart the A/B testing process.
Figure 8. Batch Size exploration with Network Functions
@@
Next, we show runtime design exploration and adaptation for two different network functions from vigor (Zaostrovnykh et al., 2019). We setup our experiment with a packet generator on one machine and a device under test (DUT) on the other machine. We use two ports for the DUT. The first port acts as the internal facing port where as the second port as the external facing port. The DUT runs executes the specific network function for every incoming packet generated by the packet generator. The network function implementation has different execution pathways for packets depending on whether they arrive on the internal port or the external port. The network function is enabled with the Iridescent specialization policy to trigger an exploration phase to find the best performing BATCH_SIZE whenever there is a change in the workload. We execute the experiment in two phases: in the first phase, packets arrive only on the external facing port and in the second phase, packets arrive only on the internal facing port. Figure 8 shows the results for the NAT and the Policer network functions. Iridescent can find the best performing configuration in each phase even if the best configuration is different.
@@
Figure 9. Optimal Fast Path size exploration with Iridescent
@@
Iridescent transparently enables exploration-based adaptation with custom specializations. We show runtime design exploration for FastClick LinearIPLookup. We execute the FastPath router on 1 machine and treat that machine as the Device Under Test and execute Pktgen to execute an open-loop workload to generate packets with destination ip addresses from a given set of IP addresses. At the 1 minute-mark, we completely switch the destination IP address set with no overlap with the initial address. The router is configured with a Iridescent specialization policy that triggers an exploration whenever it detects a large change ($\geq 25\%$) in the measured throughput. Figure 9 shows how the throughput of the router changes for the different IP sets. At the start, the specialization policy triggers an exploration as it detects a large in-flux of packets signifying a change from 0 incoming packets. First, Iridescent run a quick instrumentation phase that runs for around 100ms to find the most popular incoming destination IP addresses. After the instrumentation phase, Iridescent switches to an exploration phase to explore the size of the fast-path i.e. how many addresses should be included in the generated if-else source code. Once the exploration finishes, Iridescent switches to the best performing fast-path size and generates the specialized code which persists until the workload switch at the 1-minute mark. Due to the workload change, the specialization policy kicks in automatically, and Iridescent restarts the instrumentation and exploration phase. As shown in Figure 9, Iridescent can quickly adapt to changes in workload to generate optimal specialized code under dynamic conditions.
6.4. Iridescent but at what Cost?
@@
Compilation Cost. Table 4 shows the JIT compilation time for each of the target systems. Note that, this compilation happens off the critical path so its not a performance bottleneck. However, this cost does dictate the propagation delay, i.e., the time taken for the specialized version of the code to be available for the execution.
System JIT Compilation Time (ms)
MMulBlockBench 10 $\pm$ 1
LibLPM-FP 72 $\pm$ 9
LibLPM-NI Varies with lpm size
LibLPM-NI-FP Varies with lpm size
Network Functions 98 $\pm$ 5
TAS 340 $\pm$ 5
FastClick 11 $\pm$ 1
Table 4. JIT compilation time
@@
Figure 10. Compilation Time of LibLPM-NI (nested-if specialization) as a function of the number of elements in the LPM lookup table.
@@
Developer Cost. Table 5 shows the lines of code changes required to integrate Iridescent specializations into the target systems. For all cases, the lines of code required are fewer than 100. Implementing the custom LPM specialization only required 162 lines of code for enabling the custom user-defined LPM specialization.
System
Handler
Annotations (LoC)
Spec Policy
@@
Table 5. Lines of Code change required to integrate Iridescent
@@
Figure 11. Impact of sampling rate on instrumentation overhead for LibLPM-FP configuration for Workload B
@@
Instrumentation Overhead. To measure the overhead of adding instrumentation, we devise a microbenchmark called SimpleBench, consisting of two very simple functions, f and g. f computes the square of the input and g computes the product of its two inputs. For f, we add a general specialization point for its input a. For g, we add a range-based specialization for its second input b. For both the specialization points a and b, we first execute their respective functions, f and g, in normal mode and then use Iridescent to add instrumentation. We execute each configuration one million times. As a baseline, each function on average takes upto 8 cycles per execution in normal mode. For a, which is a general specialization point, adding instrumentation adds an overhead of around 450 to 500 cycles per execution. For b, which is a specialization point designed for more efficient instrumentation for specialization points with values in fixed ranges, adding instrumentation adds an overhead of around 1 extra cycle per operation. When possible, Iridescent users should opt to use the more efficient specialization point. However, it is expected that this might not always be possible and in most cases users would be forced to use the general specialization point, Iridescent allows users to specify a sampling rate to avoid incurring the instrumentation cost for low-latency computations. Figure 11 shows how decreasing the sampling rate for the LibLPM-FP configuration when executed with Workload B can decrease the instrumentation overhead.
@@
Specialization Guards and Failures. To measure the cost of specialization guards and failures, we use the SimpleBench microbenchmark. For function f, we use Iridescent to specialize the value of a and generate two different versions. In the first version, we disable specialization check at the entrypoint of f. We then execute this version of the function with the specialized value as input one million times. Each execution of this version takes 7 cycles. In the second version, we enable the specialization check at the entrypoint of f. If the specialization check fails, i.e. the input value of a does not match the specialized value of a, then this version of f throws an exception that is caught by Iridescent-inserted trampoline code which subsequently calls the generic version of the function. We first execute this version of the function with inputs such that the specialization guards always pass. This check adds an extra cycle per execution of the function. Next, we execute this version of the function with inputs such that the specialization guards always fail. This results in exception handling behavior which adds approximately 5000 cycles of overhead per execution. Thus, the raw cost of a specialization check is 1 extra cycle whereas that of a specialization failure is approximately 5000 cycles.
7. Related Work
Runtime Specializations. Runtime Specialization techniques (Consel and Noël, 1996; Consel et al., 2005; Grant et al., 1999; Dean et al., 1995) produce more efficient versions of the code by exploiting values and invariants that only exist at runtime. Today, these specializations manifest as optimizations either through binary rewriting (Ali and Rohou, 2017) or as optimizations in JIT compilers for interpreted languages in the form of Type Specialization (Chevalier-Boisvert et al., 2010; Ottoni, 2018; Chevalier-Boisvert and Feeley, 2015) for specializing the types of data for dynamically typed languages, or as Value Specialization for interpreted languages (de Assis Costa et al., 2014; Lima et al., 2020) where the parameter values of hot functions are converted into constants. These specializations may be optionally applied depending on the computation context (Jain and Roy, 2016; Poesia and Pereira, 2020; Flückiger et al., 2020). generic optimization and typically done automatically and transparently by the runtime based on internal cost models which may not be appropriate for specific application contexts. Moreover, these specializations are limited in scope and usually do not take the system’s end-to-end performance into consideration when applying these specializations.
Transparent Dynamic Optimization (TDO) techniques such as Dynamo (Bala et al., 2000), The Transmeta Code Morphing Software (Dehnert et al., 2003), DynamoRIO (Bruening et al., 2003), optimize code at runtime without requiring any modifications by capturing and optimizing traces (sequences of instructions) that are commonly executed. These optimization techniques are broadly applicable optimization techniques that target a broad spectrum. As a result, do not typically leverage any domain-specific, situation-specific optimizations as these techniques do not have a way of obtaining that knowledge.
Domain-Specific Specializations. In recent times, value specializations have extended beyond interpreted languages for specific use cases such as value specializations for GPU kernels (Georgakoudis et al., 2025), or through increment specializations for packet processing frameworks (Ruffy et al., 2024; Miano et al., 2022; Farshin et al., 2021). Iridescent provides a general framework for configuring, applying, exploring, and selecting such specializations at runtime.
Feedback-Driven Optimizations (FDO). FDO (Chen et al., 2016; Panchenko et al., 2019; Hölzle and Ungar, 1994; Dehnert et al., 2003; Ottoni, 2018; Selakovic et al., 2017; Stephenson and Rangan, 2021), also known as PGO (Profile-Guided Optimizations), are typically multi-run optimizations consisting of multiple executions of the program to be optimized. First the optimization collects relevant metrics from a benchmarking run and then uses these metrics to generate optimized versions of the code. FDO techniques in their current form lack flexibility as they are specialized to work well for only one workload (the workload it was trained on). FDO optimizations typically do not adapt well to change in workload or environment settings.

### v2

@@ 5.2 LibLPM
LibLPM [66] is an open-source Longest Prefix Match (LPM) library written in C with built-in support for IPv4 and IPv6 addresses. LPM is a key operation in packet routing for finding the closest matching route in routing tables for incoming packets. To integrate LibLPM with Iridescent, we create two different specializations:
LibLPM-FP
specializes the lookup function to add a fast-path specialization point. This specialization adds an exact match if-else fast-path for the top-N input IP addresses.
LibLPM-NI
@@
creates a code-generation specialization point in the LPM library. The code-generation specialization point generates a new version of the lookup function which generates a nested-if-else tree of checks consisting of prefix match checks for the incoming address for each lpm entry. The nested-if-else tree starts at the least specific matching rule (based on prefix length) and progressively checks for the most specific matching rule until it can’t find one anymore. With this code generation specialization, we embed the prefix rules directly into the codebase, forgoing expensive dynamic datastructure traversals and enabling further optimizations for individual branches.
LibLPM-NI-FP
@@
combines the previous two specializations by adding a fast-path specialization point in the generated nested-if specialized lookup function.
5.3 TAS
@@
TAS [42], TCP acceleration as a service, is a lightweight software TCP network fast-path that is optimized for common-case client-server RPCs. TAS executes common-case TCP operations in an isolated fast path that uses DPDK [25], while handling corner cases in a slow path.
@@
To integrate TAS with Iridescent, we convert its BATCH_SIZE macro to a value specialization point. In TAS, the BATCH_SIZE variable is used in three different scenarios: (i) to determine the number of packets to read from the NIC, (ii) to determine the number of packets to read from the application queues, and (iii) to determine the number of packets to read from the queue manager. For more fine-grained control, we convert each of these usage instances of the BATCH_SIZE to three separate runtime constant specialization points - (i) rx_batch, (ii) queues_batch, and (iii) qman_batch.
5.4 FastClick
@@
FastClick [7] is an extended version of the Click Modular Router [44] with Netmap and DPDK support for running the modular router in userspace. A FastClick (or Click) router is assembled from individual packet processing modules called elements. Each element implements simple router functions such as packet classification, scheduling, routing, and interacting with network devices. Users configure the router with different graphs of connected elements.
@@
To integrate FastClick with Iridescent, we modify the LinearIPLookup element of FastClick to create a Iridescent-enabled LinearIPLookup element. The original element uses a linear search algorithm to find the best matching route for an incoming packet based on the routing table. We modify the packet-processing function of the LinearIPLookup element to add a fast-path specialization point. The target of this specialization point is the internal lookup function of the LinearIPLookup element.
5.5 Network Functions
@@
Network Functions is a suite of the network functions that include DPDK and ebpf implementations of common network functions such as a NAT, Router, policer, among others. These functions have been developed and used for evaluation by various research projects such as Pix [36] and Vigor [88]. We extract the implementations from the Pix artifact [35].
@@
6 Evaluation
In this section, we evaluate how much Iridescent improves performance of systems by enabling online specializations. We show how we can use online specialization with Iridescent for three use-cases: (i) enabling compile-time optimizations at runtime; (ii) enabling incremental specializations at runtime; (iii) design exploration at runtime. We answer the following questions:
  • Do Iridescent-enabled optimizations improve performance? (§ 6.2)
  • Can Iridescent find an optimal specialization through exploration at runtime under dynamic conditions? (§ 6.3)
  • What is the overhead of using Iridescent? (§ 6.4)
6.1 Experimental Setup
Testbed.
For our experiments, unless otherwise stated, we configure two identical machines as client and server. They are directly connected with a pair of 100 Gbps Mellanox ConnectX-5 Ethernet adapters. Each machine has two Intel Xeon Gold 6152 processors at 2.1 GHz, each with 22 cores for a total of 44 cores and 187 GB of RAM per machine. We run Linux kernel 5.15 with Debian 11.
6.2 Compiler Optimizations at Runtime
@@
Table 4: Impact of turning block size, $B$, as compile-time constant at runtime for N=64 for different hardwares (Table 1)
Figure 3: Average cycles required to perform a table lookup with varying table sizes under two different workloads.
Compile-time optimizations often produce more performant code. These compile time optimizations include constant propagation, loop unrolling, dead-code elimination, and use of vectorization instructions. With Iridescent, we can enable compile time optimizations using runtime data.
Iridescent enables cascading compiler optimizations.
First, we show the potential benefits of enabling compiler optimizations at runtime for the MMulBlockBench microbenchmark in Table 1. Specifically, for the workload $N=64$, we compare the performance of keeping the optimal block size, $B$, as a runtime parameter to that of converting it into a constant at runtime using Iridescent, thus enabling downstream compiler optimizations. For both configurations, we execute the function a fixed number of times and measure the number of cycles spent per function execution. Table 4 shows the benefit of converting $B$ into a compile-time constant for different machines. For each machines, we get at least 50% reduction in cycles, and greater than 240% reduction in cycles for four of the five operating conditions. This improvement is a direct impact of Iridescent’s specialization allowing different compiler optimizations to cascade and combine together to produce a more efficient version of the code. In this specific case, the specialization of $B$ allows the JIT compiler to first easily unroll the loops in the matrix multiplication function and then leverage optimized vector instructions.
Figure 4: FastPath router throughput with Iridescent specialized LinearIPLookup element for different routing table sizes at 100% fast-path hit rate.
Figure 5: FastPath router throughput with Iridescent specialized LinearIPLookup element for varying fast-path hit rates.
Iridescent incorporates existing specializations.
@@
Next, we show the benefits of custom compile-time optimizations for LibLPM. For this experiment, we set up eleven different configurations resulting from a combination of five different LPM table sizes and two different workloads - Workload A and Workload B. In Workload A, the incoming IP address is an IP address that matches a very specific prefix entry in the routing table. In Workload B, the incoming IP address is an IP address that only matches the LPM prefix entry of prefix length 0. Thus, Workload A and Workload B cover the best and worst cases respectively for the LPM lookup function. We execute these workloads with seven specializations:
  1. (i)
    Baseline: no Iridescent-enabled specialization.
  2. (ii)
    Instrumented: Iridescent-enabled specialization capturing the most popular inputs (10% sampling); data used to guide the fast path specializations.
  3. (iii)
    LibLPM-FP-A: fast-path specialization of size 1 specialized for Workload A.
  4. (iv)
    LibLPM-FP-B: fast-path specialization of size 1 specialized for Workload B.
  5. (v)
    LibLPM-NI: nested-if code generation specialization.
  6. (vi)
    LibLPM-NI-FP-A: Combine LibLPM-NI and -FP-A.
  7. (vii)
    LibLPM-NI-FP-B: Combine LibLPM-NI and -FP-B.
Figure 3 shows average cycles per LPM lookup lookup function for all different combinations of configurations and specializations. For Workload A, all three specializations, LibLPM-FP-A, LibLPM-NI, LibLPM-NI-FP-A, achieve up to $6\times$ reduction in cycles per execution. However, unlike the LibLPM-FP and LibLPM-NI-FP-A specialization, this LibLPM-NI specialization does not require an additional instrumentation phase to sample executions to find the optimal candidates for the fast path. LibLPM-NI improves performance across the board for all workloads by up to $6\times$. Despite this performance boost, LibLPM-NI is not the best specialization for Workload B. For Workload B, the best optimizations are the *-FP-B specializations as they improve by up to $30\times$ as all instantiations of Workload B hit the fast-path.
6.3 Runtime Design Exploration
Figure 6: Automatic exploration and specialization of different configurations for TAS.
Iridescent enables runtime design exploration.
@@ 6.3 Runtime Design Exploration
We show runtime design exploration for TAS [42]. We experiment in a typical server-client setting, with the server providing an echo service. The client is multi-threaded and executes an open-loop workload of 64 byte packets. Both the server and client are TAS-enabled. In our experiment, we modify the server-side TAS and measure the throughput of the server side TAS as millions of packets processed per second. With Iridescent we explore different values of the rx_batch in the TAS source code on the server side. Figure 6 shows Iridescent automatically determining and then selecting the best-configuration on the server side by exploring the different values of the batch size for different code locations.
Figure 7: Automatic exploration and specialization of MMulBlockBench for two different workloads.
Iridescent automatically adapts to changing workloads.
@@
To show that Iridescent can automatically adapt to workloads, we first use the MMulBlockBench microbenchmark. We execute a Iridescent-enabled version of the function and a non-Iridescent enabled baseline. In the baseline, the block size $B$ is a runtime parameter. In the Iridescent-enabled version, Iridescent explores different values of $B$ by specializing the value in the code and converting $B$ into a compile-time constant. We execute each version with two different workloads in succession, each lasting one minute. We measure the overall throughput (executions/s) for both configurations. Figure 7 shows how Iridescent explores the different block sizes and finds the most performant configuration compared to the baseline. Moreover, Iridescent can automatically detect the workload change based on the drop in throughput and restart the A/B testing process.
Figure 8: Batch size exploration with Network Functions
@@
Next, we show runtime design exploration and adaptation for two different network functions, vignat and vigpol, from vigor [88]. We setup our experiment with a packet generator on one machine and a device under test (DUT) on the other machine. We use two ports for the DUT. The first port acts as the internal facing port whereas the second port as the external facing port. The DUT executes the specific network function for every incoming packet generated by the packet generator. The network function implementation has different execution pathways for packets depending on whether they arrive on the internal port or the external port. The network function is enabled with the Iridescent specialization policy to trigger an exploration phase to find the best performing BATCH_SIZE whenever there is a change in the workload. We execute the experiment in two phases: in the first phase, packets arrive only on the external facing port and in the second phase, packets arrive only on the internal facing port. Figure 8 shows the results for the NAT and the Policer network functions. Iridescent finds the best performing configuration in each phase with different optimal values.
@@
Figure 9: Optimal fast path size exploration with Iridescent.
Iridescent transparently enables exploration-based adaptation with custom specializations.
@@
We show runtime design exploration for FastClick LinearIPLookup. We run the FastPath router on one machine, treat that machine as the Device Under Test and run Pktgen to execute an open-loop workload generating packets with destination ip addresses from a given set of IP addresses. At the 1 minute-mark, we completely switch the destination IP address set with no overlap with the initial address. The router is configured with a Iridescent specialization policy that triggers an exploration whenever it detects a large change ($\geq 25\%$) in measured throughput. Figure 9 shows how the throughput of the router changes for the different IP sets. At the start, the specialization policy triggers an exploration as it detects a large in-flux of packets, a significant change from 0 incoming packets. First, Iridescent runs a quick instrumentation phase for around 100ms to find the most popular incoming destination IP addresses. After the instrumentation phase, Iridescent switches to an exploration phase to explore the size of the fast-path, i.e., how many addresses should be included in the generated if-else source code. Once the exploration finishes, Iridescent switches to the best performing fast-path size and generates the specialized code which persists until the workload switch at the 1-minute mark. Due to the workload change, the specialization policy kicks in automatically, and Iridescent restarts the instrumentation and exploration phase. As shown in Figure 9, Iridescent can quickly adapt to changes in workload to generate optimal specialized code under dynamic conditions.
Iridescent can explore large specialization spaces.
We explore the large specialization space of CloverLeaf with OPPerTune’s exploration algorithm integrated with Iridescent. We limit the number of configurations to be explored to 20 out of a total of $7^{24}$ total configurations. We observe that, Iridescent can find an configuration that reduces the number of cycles spent per iteration of the main CloverLeaf loop by $3\times$ on average. By leveraging state-of-the-art auto-tuning exploration strategies, Iridescent can efficiently find an optimal configuration from large specialization spaces.
6.4 Iridescent but at what Cost?
Compilation cost.
@@ 6.4 Iridescent but at what Cost?
Table 5 shows the JIT compilation time for each of the target systems. Note that, this compilation happens off the critical path so it is not a performance bottleneck. However, this cost does dictate the propagation delay, i.e., the time taken for the specialized version of the code to be available for the execution.
System Time (ms) Size (KB)
MMulBlockBench 10 $\pm$ 1 10
LibLPM-FP 72 $\pm$ 9 80
Network Functions 98 $\pm$ 5 67
TAS 340 $\pm$ 5 360
FastClick 11 $\pm$ 1 9.2
Table 5: JIT compilation time and corresponding size
@@
Figure 10: Compilation time of LibLPM-NI (nested-if specialization) as a function of the number of elements in the LPM lookup table.
@@
Developer cost.
Table 6 shows the necessary lines of code changes to integrate Iridescent specializations into the target systems. For all cases, fewer than 100 lines of code change. Implementing the custom LPM specialization only required 162 lines of code for enabling the custom user-defined LPM specialization.
System Handler Annotations (LoC) Spec Policy (LoC)
@@
Table 6: Lines of Code change required to integrate Iridescent.
@@
Figure 11: Impact of sampling rate on instrumentation overhead for LibLPM-FP configuration for Workload B.
Instrumentation overhead.
@@
To measure the overhead of adding instrumentation, we devise a microbenchmark called SimpleBench, consisting of two very simple functions, f and g. f computes the square of the input and g computes the product of its two inputs. For f, we add a general specialization point for its input a. For g, we add a range-based specialization for its second input b. For both the specialization points a and b, we first execute their respective functions, f and g, in normal mode and then use Iridescent to add instrumentation. We execute each configuration one million times. As a baseline, each function on average takes up to 8 cycles per execution in normal mode. For a, which is a general specialization point, adding instrumentation adds an overhead of around 450 to 500 cycles per execution. For b, which is a specialization point designed for more efficient instrumentation for specialization points with values in fixed ranges, adding instrumentation adds an overhead of around 1 extra cycle per operation. When possible, Iridescent users should opt to use the more efficient specialization point. However, as this ma not always be possible and users may be forced to use the general specialization point, Iridescent allows users to specify a sampling rate to avoid incurring the instrumentation cost for low-latency computations. Figure 11 shows how decreasing the sampling rate for the LibLPM-FP Workload B configuration can decrease the instrumentation overhead.
Specialization guards and failures.
@@
To measure the cost of specialization guards and failures, we use the SimpleBench microbenchmark. For function f, we use Iridescent to specialize the value of a and generate two different versions. In the first version, we disable the specialization check at the entrypoint of f. We then execute this version of the function with the specialized value as input one million times. Each execution of this version takes 7 cycles. In the second version, we enable the specialization check at the entrypoint of f. If the specialization check fails, i.e. the input value of a does not match the specialized value of a, then this version of f throws an exception that is caught by Iridescent-inserted trampoline code which subsequently calls the generic version of the function. We first execute this version of the function with inputs such that the specialization guards always pass. This check adds an extra cycle per execution of the function. Next, we execute this version of the function with inputs such that the specialization guards always fail. This results in exception handling behavior which adds approximately 5000 cycles of overhead per execution. Thus, the raw cost of a specialization check is 1 extra cycle whereas that of a specialization failure is approximately 5000 cycles.
7 Related Work
7.1 Offline Automatic Specializations
We classify offline automatic specialization techniques in four categories. (i) Compile pass selection techniques automatically select the optimal sequence of compiler optimization passes through auto-tuning methods for trying out different combinations [62, 58, 59, 58], (ii) Specialized compiler optimization passes that can generate specialized code for specific hardware features like vectorization [56] or prefetching [73], or for other specific optimizations [47, 52, 69, 2], (iii) Feedback-Driven Optimizations (FDO) [13, 61, 34, 23, 57, 71, 79, 67, 87, 33, 90, 38, 89, 33], also known as PGO (Profile-Guided Optimizations), are compiler optimizations consisting of multiple executions of the program to be optimized by collecting relevant metrics from a benchmarking run and then using these metrics to generate optimized versions. FDO techniques in their current form lack flexibility as they are specialized to work well for only one workload (the workload it was trained on), (iv) Specialized Code Tuning approaches either use trial runs to tune specific parameters in the code [46, 64] or apply specialized optimizations to different pieces of the code [81, 64]. Unlike Iridescent, existing offline automatic optimizations cannot adapt to changes in workload or environmental settings dynamically. Consequently, these techniques are sub-optimal for the conditions they were not trained on.
7.2 Online Automatic Optimizations
We classify existing online automatic optimizations in four categories — (i) runtime code selection, (ii) configuration optimizations, (iii) generic code specializations, (iv) domain-specific specializations.
Runtime code selection.
In this optimization technique, developers pre-generate multiple different versions of the code at compile-time and then use measured runtime statistics and offline-trained models to switch between the versions [40, 76, 78, 20]. These techniques are limited as they must anticipate and generate all possible versions a priori. Consequently, in contrast to Iridescent, these techniques can not fully leverage runtime information for specialization.
Config optimizations.